OpenStudy (anonymous):

What is the simplified expression for 5 to the power of 4 multiplied by 5 to the power of 3 over 5 to the power of 5?

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

\(\sf \Large a^n \times b^n \times c^n=(a\times b\times c)^n\)

OpenStudy (abhisar):

Can you do it now?

OpenStudy (anonymous):

\[a^m \times a^n=a^{m+n}\] \[\frac{a^m}{a^n}=a^{m-n}\] \[\frac{a^l \times a^m}{a^n}=a^{l+m-n}\]

OpenStudy (abhisar):

Oh yes, it says 5 to the power 4 which means \(\sf 5^4\), so you will have to use \(\sf a^x \times a^y \times a^z = a^{x+y+z}\)